# Map-level baryonification: Efficient modelling of higher-order correlations in the weak lensing and thermal Sunyaev-Zeldovich fields

### v1

@@ Abstract
Semi-analytic methods can generate baryon-corrected fields from N-body simulations (“baryonification”) and are rapidly becoming a ubiquitous tool in modeling structure formation on non-linear scales. We extend this formalism to consistently model the weak lensing and thermal Sunyaev-Zeldovich (tSZ) fields directly on the full-sky, with an emphasis on higher-order correlations. We use the auto- and cross- $N$th-order moments, with $N\in\{2,3,4\}$, as a summary statistic of the lensing and tSZ fields, and show that our model can jointly fit these statistics measured in IllustrisTNG to within measurement uncertainties, for scales above $\gtrsim 1\,{\rm Mpc}$ and across multiple redshifts. The model predictions change only minimally when including additional information from secondary halo properties, such as halo concentration and ellipticity. Each individual moment is dependent on halos of different mass ranges and has different sensitivities to the model parameters. A simulation-based forecast on the Ulagam suite shows that the combination of all moments, measured from current and upcoming lensing and tSZ surveys, can jointly constrain cosmology and baryons to high precision. The lensing and tSZ field are sensitive to different combinations of the baryonification parameters, with degeneracy directions that are often orthogonal, and the combination of the two fields leads to significantly better constraints on both cosmology and astrophysics. Our pipeline for map-level baryonification is publicly available at https://github.com/DhayaaAnbajagane/Baryonification.
@@ 1 Introduction
Given the current challenges in modelling baryons in simulations, it is advantageous to instead use a phenomenological model. Such a model does not require specifying any sub-grid physics and instead requires inputs that are observables, or are quantities directly obtained from observables, of the different baryonic matter components. Baryonification is a method through which N-body simulations — which evolve collisionless matter (dark matter) under gravity and accurately capture all non-linear evolution — are modified through phenomonological models such that the new density field mimics one that has baryonic imprints (Schneider & Teyssier 2015; Schneider et al. 2019). This method has been used extensively to model the matter power spectrum with baryonic imprints, and has been shown to have the flexibility needed to capture the behaviors ranging across many different simulations (Schneider et al. 2019; Giri & Schneider 2021; Aricò et al. 2021b).
@@
Pandey et al. (2024, henceforth, P24) show that the existing baryonification model can be extended, with simple modifications, to also predict the gas pressure profiles (and gas number density profile) in addition to the total matter density profiles. To et al. (2024, see their Figure 1) have further demonstrated that this model can consistently predict the tSZ–halo mass scaling relation alongside the density two-point correlations. Thus, the original baryonification model of Schneider et al. (2019, henceforth, S19) can be built on to now also predict the tSZ observables (as well as other observables that trace the gas number density, such as the kinematic SZ effect or the X-ray flux). While P24 focus on using the halo model approach to predict two-point measurements of weak lensing and the tSZ, the model remains to be explored and validated at the map-level for different summary statistics that combine information from the lensing and tSZ fields.
@@
A number of appendices provide further details on the methods we used: the lensing and tSZ forward modelling approach in Appendix A, the sensitivity of the predictions to additional model and methodology choices in Appendix B and C, the redshift evolution of the baryonification predictions in Appendix D, and finally, some salient computation details on the method, include the runtimes, in Appendix E.
@@ 2 Baryonification
A number of salient computational details about the method are described in Appendix E, including the characteristic runtimes of the pipeline. The baryonified lensing and tSZ maps for a simulation [^2] Our fiducial simulation contains around 2 million halos, and we baryonify around 90 density shells from $0<z\lesssim 3.5$, each shell a HealPix map of NSIDE = 1024. See Section 3.3 for more details. can be generated on the order of minutes on an Intel broadwell chip with 40 cores; see Appendix E for more details.
@@ 2.1 Dark matter baryon (DMB) halo model
The DMB halo model is a sum over multiple different components. Each component is parameterize in a different way, with the parameterization informed by prior work in simulations and/or observations. We now describe each component below:
@@
Here, the two power-law exponents are given by $\tau_{\rm cga}=\tau+\tau_{\rm\delta}$ and $\eta_{\rm cga}=\eta+\eta_{\rm\delta}$, where $\tau$ and $\eta$ correspond to the scaling of the total stellar fraction, and $\tau_{\delta}$ and $\eta_{\delta}$ are a correction to convert the total stellar fraction scaling to the central galaxy stellar fraction scaling. The factor of 2 ensures that for $M_{\rm 200c}=M_{1}$, we have $f_{\rm cga}=A$. The original baryonification model of S19 used a different version of Equation (9), which included $\eta$ and $\eta_{\delta}$ but did not have $\tau$ and $\tau_{\delta}$. These latter parameters control the slope of the stellar-mass halo-mass relation for low-mass halos and is included in this work. In practice, the $\tau$ and $\tau_{\delta}$ parameters have no impact in our work as the signal for both weak lensing and tSZ is generated by halos that are above the $M_{1}=3\times 10^{11}\,{\rm M}_{\odot}$ mass scale (see Table 1 for the fiducial values of different parameters), and such halos are sensitive only to $\eta,\eta_{\delta}$ and not $\tau,\tau_{\delta}$. However, we include this parameter here as it is a simple extension that generalizes the model for smaller mass halos.
@@
with the definition $f_{\rm CLM}=(\Omega_{\rm m}-\Omega_{b})/\Omega_{\rm m}-(f_{\rm star}-f_{\rm cga})$. The latter two fractions are defined in Equation (9) and (15), and their difference $f_{\rm star}-f_{\rm cga}$ corresponds to the mass fraction of stars in satellite galaxies. This quantity, $f_{\rm CLM}$, corresponds to the fraction of collisionless matter in the total matter distribution.
@@ 2.2 Matter density field
This choice of projection scale can also be motivated by taking the limit $L_{\rm p}\rightarrow\infty$, where $M_{\rm DMB,p}\approx M_{\rm DMO,p}$ since the large-scale matter distribution (which is much less affected by baryons than the small-scale distribution, and is modelled by the two-halo term in our work) dominates the mass in the pixel. In this limit, the displacement function is simply $\Delta d(r_{\rm p})\approx 0$. By setting $L_{\rm p}$ as the integration distance, our model predictions are formally consistent in these extreme limits. Appendix B shows the changes in the predicted displacement due to variations in the assumed projection scale.
@@ 2.3 Thermal Sunyaev-Zeldovich (tSZ) effect
Aricò & Angulo (2024) also introduce an extension of their baryonification method (Aricò et al. 2021b) to include a tSZ field consistent with the baryon-corrected density field. They first assign “synthetic” gas and star particles to their simulations using the DM particle dataset. The pressure profile is then modelled using the prescription of Komatsu & Seljak (2002). That prescription connects the gas pressure and density by assuming the gas is a polytrope ($P\propto\rho^{\Gamma}$), which is allowed under the condition of hydrostatic equilibrium, and by assuming the total matter distribution follows an NFW profile. Note that $\Gamma$ is furthermore assumed to be independent of scale, whereas it has been found to be a function of radius in both simulations (e.g., Battaglia et al. 2012, see their Figure 4) and observations (Capelo et al. 2012). Finally, a particle at a given radial distance, $r$, is assigned a value given by the pressure profile evaluated at that distance.
In comparison, our work predicts the pressure field directly on the sky, avoiding the use of particles to reduce computational cost and relying instead on the halo positions. Our pressure profile is obtained using the assumptions of hydrostatic equilibrium as well, but we do not assume an NFW profile for the total matter distribution. Instead, we directly use $M_{\rm dmb}(r)$ predicted by the baryonification model; see Equation (22). Our model also does not assume the gas is a polytrope or that it has a scale-independent polytropic index. This allows our predictions to be more easily generalized to both the outskirts of halos and to lower-mass halos.
@@ 2.5 Model parameterization
$M_{\rm c}$ $10^{14.5}\,{\rm M}_{\odot}$ [$10^{10}$, $10^{18}\,{\rm M}_{\odot}$] Gas Mass scale at which half the gas mass has been ejected from the halo. Determines $\beta$
@@ 3.1 IllustrisTNG simulations
Each simulation in the suite also has a DMO counterpart, that shares the exact same initial conditions. The pair of simulations can be used to make measurements whose uncertainties are cosmic variance suppressed (see Section 3.5 for details) and therefore allow more precise validations than otherwise possible. There also exist other, public simulation suites that span much large volumes than IllustrisTNG — such as Magneticum (Hirschmann et al. 2014) and Bahamas (McCarthy et al. 2017) — and would therefore be better suited for our tests of the tSZ field and lensing field cross-correlations. However, unlike the case with IllustrisTNG, the quantities we require from these simulations (the particle snapshots) are not publicly accessible.
@@ 3.3 Ulagam simulations
Modelling the cross-correlation of the lensing and tSZ field through these simulations incurs one caveat. Pandey et al. (2022, see their Figure 2) compute the mass-dependence of the lensing and tSZ real-space, two-point cross-correlation function. For the smallest scales we measure ($\arcmin\theta\approx 3\arcmin$) they show the peak contributing halo mass is $M_{\rm 200c}\approx 10^{14}\,{\rm M}_{\odot}$ but halos below that mass-scale also contribute a significant amount to the total signal. Our map will therefore underestimate this signal since we only paint pressure profiles around halos with a minimum mass of $10^{14}\,{\rm M}_{\odot}$. Higher-order correlations push the peak contributing halo mass to higher masses, therefore making any missing contribution from lower mass halos less relevant. Regardless, we stress that the missing lower mass halos in our analysis only serve to reduce the amplitude of the simulated signal and therefore will weaken our fisher information. We do not risk quoting artificially tighter bounds due to this caveat. In summary, the Ulagam suite still has sufficient, though not ideal, characteristics for performing our joint analysis of lensing and tSZ fields.
@@ 3.5 Estimation of measurement uncertainty and model best fit
Equation (34) uses the tSZ field of the hydrodynamical simulation and modifies it in Fourier space to now correspond to the density field of the DMO simulation. This field is then used in estimating $R$ for moments of the tSZ field.
@@ 4.1.1 Comparing Particle and Shell baryonification
For our tSZ profile painting method, we can interchangeably use the 3D and 2D approach as the difference is simply in what step the line-of-sight integration is done. In the first method, we paint profiles on a 3D grid (we do not use particle information when painting profiles) and then integrate along line-of-sight by collapsing the map along one direction. In the second, the integration is done within the definition of the projected, 2D profile, before any painting happens. Thus, the tSZ field does not suffer accuracy drops between using 3D fields and 2D fields.
@@
Once we have made the map for a given model, we take the fractional difference of the moments measured on the two maps from particle/shell baryonification, and then compute the average absolute ratio over all 32 models. Figure 3 shows this average difference for the moments of the density field. In practice, the offsets are within the 1-2% level, which is a negligible error. The error in the third moments increases at around the pixel-scale of $\approx 1\,{\rm Mpc}$, and the same for the fourth moments increases below $\lesssim 2\,{\rm Mpc}$. These differences are still within the cosmic variance-suppressed uncertainties of the measurement (see gray bands in Figure 8).
Figure 3: The (absolute) fractional difference in the baryonification predictions, averaged over 32 parameter choices, between the 2D shell method and the 3D particle method (see Section 2.2 for details on both methods). The 32 models are a Sobol sequence drawn from a prior defined in Table 1. The bias is within $2\%$ across all scales and increases for higher-order statistics for scales below $\lesssim 2\,{\rm Mpc}$. We deem this a negligible systematic; see text for more details.

### v2

@@ Abstract
Semi-analytic methods can generate baryon-corrected fields from N-body simulations (“baryonification”) and are rapidly becoming a ubiquitous tool in modeling structure formation on non-linear scales. We extend this formalism to consistently model the weak lensing and thermal Sunyaev-Zeldovich (tSZ) fields directly on the full-sky, with an emphasis on higher-order correlations. We use the auto- and cross- $N$th-order moments, with $N\in\{2,3,4\}$, as a summary statistic of the lensing and tSZ fields, and show that our model can jointly fit these statistics measured in IllustrisTNG to within measurement uncertainties, for scales above $\gtrsim 1\,{\rm Mpc}$ and across multiple redshifts. The model predictions change only minimally when including additional information from secondary halo properties, such as halo concentration and ellipticity. Each individual moment is dependent on halos of different mass ranges and has different sensitivities to the model parameters. A simulation-based forecast on the Ulagam suite shows that the combination of all moments, measured from current and upcoming lensing and tSZ surveys, can jointly constrain cosmology and baryons to high precision. The lensing and tSZ field are sensitive to different combinations of the baryonification parameters, with degeneracy directions that are often orthogonal, and the combination of the two fields leads to significantly better constraints on both cosmology and astrophysics. Our pipeline for map-level baryonification is publicly available at https://github.com/DhayaaAnbajagane/BaryonForge.
@@ 1 Introduction
Given the current challenges in modelling baryons in simulations, it is advantageous to instead use a phenomenological model. Such a model does not require specifying any sub-grid physics and instead requires inputs that are observables, or are quantities directly obtained from observables, of the different baryonic matter components. Baryonification is a method through which N-body simulations — which evolve collisionless matter (dark matter) under gravity and accurately capture all non-linear evolution — are modified through phenomonological models such that the new density field mimics one that has baryonic imprints (Schneider & Teyssier 2015; Schneider et al. 2019). This method has been used extensively to model the matter power spectrum with baryonic imprints, and has been shown to have the flexibility needed to capture the behaviors ranging across many different simulations (Schneider et al. 2019; Giri & Schneider 2021; Aricò et al. 2021b). It has already been used to analyse 2-point correlation functions measured in widefield surveys (Chen et al. 2023; Aricò et al. 2023; Bigwood et al. 2024).
@@
Pandey et al. (2024, henceforth, P24) show that the halo profiles used in the existing baryonification model can be extended, with simple modifications, to also predict the gas pressure profiles (and gas number density profile) in addition to the total matter density profiles. To et al. (2024, see their Figure 1) have further demonstrated that this model can consistently predict the tSZ–halo mass scaling relation alongside the density two-point correlations. Thus, the original baryonification model of Schneider et al. (2019, henceforth, S19) can be built on to now also predict the tSZ observables (as well as other observables that trace the gas number density, such as the kinematic SZ effect or the X-ray flux). While P24 focus on using the halo model approach to predict two-point measurements of weak lensing and the tSZ, the model remains to be explored and validated at the map-level for different summary statistics that combine information from the lensing and tSZ fields.
@@
A number of appendices provide further details on the methods we used: the lensing and tSZ forward modelling approach in Appendix A, the sensitivity of the predictions to additional model and methodology choices in Appendix B and C, the redshift evolution of the baryonification predictions in Appendix D, the dependence between Fisher information and baryonification model-complexity in Appendix E, and finally, some salient computation details on the method, include the runtimes, in Appendix F.
@@ 2 Baryonification
A number of salient computational details about the method are described in Appendix F, including the characteristic runtimes of the pipeline. The baryonified lensing and tSZ maps for a simulation [^2] Our fiducial simulation contains around 2 million halos, and we baryonify around 90 density shells from $0<z\lesssim 3.5$, each shell a HealPix map of NSIDE = 1024. See Section 3.3 for more details. can be generated on the order of minutes on an Intel broadwell chip with 40 cores; see Appendix F for more details.
@@ 2.1 Dark matter baryon (DMB) halo model
The DMB halo model is a sum over multiple different components. Each component is parameterized in a different way, with the parameterization informed by prior work in simulations and/or observations. We now describe each component below:
@@
Here, the two power-law exponents are given by $\tau_{\rm cga}=\tau+\tau_{\rm\delta}$ and $\eta_{\rm cga}=\eta+\eta_{\rm\delta}$, where $\tau$ and $\eta$ correspond to the scaling of the total stellar fraction, and $\tau_{\delta}$ and $\eta_{\delta}$ are a correction to convert the total stellar fraction scaling to the central galaxy stellar fraction scaling. The factor of 2 ensures that for $M_{\rm 200c}=M_{1}$, we have $f_{\rm cga}=A$. The original baryonification model of S19 used a different version of Equation (9), which included $\eta$ and $\eta_{\delta}$ but did not have $\tau$ and $\tau_{\delta}$. These latter parameters control the slope of the stellar-mass halo-mass relation for low-mass halos and is included in this work. [^3] S19 note that analyses of weak lensing only use halos above $M_{\rm 200c}>10^{12}\,{\rm M}_{\odot}$, for which their single power-law model is accurate and adequate. We extend this to the double power-law model traditionally used in observational and theoretical analyses of galaxies (e.g., Behroozi et al. 2019, see their Figure 9). This extension allows the baryonification model to be used consistently with halos of lower masses. In practice, the $\tau$ and $\tau_{\delta}$ parameters have no impact in our work as the signal for both weak lensing and tSZ is generated by halos that are above the $M_{1}=3\times 10^{11}\,{\rm M}_{\odot}$ mass scale (see Table 1 for the fiducial values of different parameters), and such halos are sensitive only to $\eta,\eta_{\delta}$ and not $\tau,\tau_{\delta}$. However, we include this parameter here as it is a simple extension that generalizes the model for smaller mass halos.
@@
with the definition $f_{\rm CLM}=(\Omega_{\rm m}-\Omega_{b})/\Omega_{\rm m}+(f_{\rm star}-f_{\rm cga})$. The latter two fractions are defined in Equation (9) and (15), and their difference $f_{\rm star}-f_{\rm cga}$ corresponds to the mass fraction of stars in satellite galaxies. This quantity, $f_{\rm CLM}$, corresponds to the fraction of collisionless matter in the total matter distribution.
@@ 2.2 Matter density field
This choice of projection scale can also be motivated by taking the limit $L_{\rm p}\rightarrow\infty$, where $M_{\rm DMB,p}\approx M_{\rm DMO,p}$ since the large-scale matter distribution (which is much less affected by baryons than the small-scale distribution, and is modelled by the two-halo term in our work) dominates the mass in the pixel. In this limit, the displacement function is simply $\Delta d(r_{\rm p})\approx 0$. By setting $L_{\rm p}$ as the integration distance, our model predictions are formally consistent in these extreme limits. Appendix B shows the changes in the predicted displacement due to variations in the assumed projection scale. We note that the projection integral in Equation (25) assumes the halo is spherically symmetric. We test the anisotropic baryonification corrections in Section 4.1.3, but the baryonification model in those cases will still use the integral in Equation (25).
@@
We have verified that our particle baryonification pipeline, operating on the IllustrisTNG DMO simulation, can reproduce the 3D power spectrum results from Figure 2 of S19. We have also checked — using the shell baryonification pipeline — that the angular power spectrum of the baryon-corrected density fields shows the same scale-dependent variation with model parameters as in that Figure; for example, using $\theta_{\rm ej}=2$ leads to an excess of power on intermediate scales. [^6] For brevity, we do not include plots of these validation tests, but they are found (and can be easily reproduced by any user) in Jupyter notebooks in our public repository; see the Data Availability Section below for more details on accessing the repository.
@@ 2.3 Thermal Sunyaev-Zeldovich (tSZ) effect
Aricò & Angulo (2024) also introduce an extension of their baryonification method (Aricò et al. 2021b) to include a tSZ field consistent with the baryon-corrected density field. They first assign “synthetic” gas and star particles to their simulations using the DM particle dataset. The pressure profile is then modelled using the prescription of Komatsu & Seljak (2002). That prescription connects the gas pressure and density by assuming the gas is a polytrope ($P\propto\rho^{\Gamma}$), which is allowed under the condition of hydrostatic equilibrium, and then by assuming the total matter distribution follows an NFW-like profile (Komatsu & Seljak 2001, see their Equation 7). Note that $\Gamma$ is furthermore assumed to be independent of scale, whereas it has been found to be a function of radius in both simulations (e.g., Battaglia et al. 2012, see their Figure 4) and observations (Capelo et al. 2012); it does, however, depend explicitly on halo concentration and therefore halo mass. Finally, a particle at a given radial distance, $r$, is assigned a value given by the pressure profile evaluated at that distance.
In comparison, our work predicts the pressure field directly on the sky, avoiding the use of particles to reduce computational cost and relying instead on the halo positions. Our pressure profile is obtained using the assumptions of hydrostatic equilibrium as well, but we do not assume an NFW profile for the total matter distribution. Instead, we directly use $M_{\rm dmb}(r)$ predicted by the baryonification model; see Equation (22). Our model also does not assume the gas is a polytrope or that it has a scale-independent polytropic index. This allows our predictions to be more easily generalized to both the outskirts of halos and to lower-mass halos. Our profile painting method can capture anisotropic features in the field via the use of elliptical profiles (see Section 4.1.3), but it does not capture more fine-grained morphological features, as would be possible in the particle-based method of Aricò & Angulo (2024).
@@ 2.5 Model parameterization
$M_{\rm c}$ $10^{14.5}\,{\rm M}_{\odot}$ [$10^{10}$, $10^{18}\,{\rm M}_{\odot}$] Gas Mass-scale where the characteristic gas slope has value, $\beta=1.5$
@@ 3.1 IllustrisTNG simulations
Each simulation in the suite also has a DMO counterpart, that shares the exact same initial conditions. The pair of simulations can be used to make measurements whose uncertainties are cosmic variance suppressed (see Section 3.5 for details) and therefore allow more precise validations than otherwise possible. We use the public data release, as introduced in Nelson et al. (2019). There also exist other, public simulation suites that span much large volumes than IllustrisTNG — such as Magneticum (Hirschmann et al. 2014) and Bahamas (McCarthy et al. 2017) — and would therefore be better suited for our tests of the tSZ field and lensing field cross-correlations. These simulations also span a wide range of baryonic feedback prescriptions, some of which are stronger compared to the relatively weaker feedback in IllustrisTNG. However, unlike the case with IllustrisTNG, the quantities we require from these simulations (the particle snapshots) are not publicly accessible. For these reasons, we use IllustrisTNG for our initial validation here, but highlight the need for a more precise validation — using a larger simulation — before employing this model on survey data.
@@ 3.3 Ulagam simulations
Modelling the cross-correlation of the lensing and tSZ field through these simulations incurs one caveat. Pandey et al. (2022, see their Figure 2) compute the mass-dependence of the lensing and tSZ real-space, two-point cross-correlation function. For the smallest scales we measure ($\arcmin\theta\approx 3\arcmin$) they show the peak contributing halo mass is $M_{\rm 200c}\approx 10^{14}\,{\rm M}_{\odot}$ but halos below that mass-scale also contribute a significant amount to the total signal. Our map will therefore underestimate this signal since we only paint pressure profiles around halos with a minimum mass of $10^{14}\,{\rm M}_{\odot}$. Higher-order correlations push the peak contributing halo mass to higher masses, therefore making any missing contribution from lower mass halos less relevant. Regardless, we stress that the absence of lower mass halos in our analysis will reduce the impact of baryons on our simulated signal (relative to their impact when using all halos) and therefore will weaken our Fisher information on baryonic processes. We do not risk quoting artificially tighter bounds due to this caveat. In summary, the Ulagam suite still has sufficient, though not ideal, characteristics for performing our joint analysis of lensing and tSZ fields.
@@ 3.5 Estimation of measurement uncertainty and model best fit
Equation (34) uses the tSZ field of the hydrodynamical simulation and modifies it in Fourier space to now correspond to the density field of the DMO simulation. This field is then used in estimating $R$ for moments of the tSZ field. This new “artificial” tSZ field accounts for differences in the exact realizations of the hydrodynamical and DMO simulations, which requires information about fourier-mode phases, and also the differences in the clustering of matter on small scales, which requires information about the fourier-mode amplitudes. Note that there are many other reasonable ways to estimate a “artificial” tSZ field, such as the combination $y_{\rm dmo}={\rm iFFT}({\rm FFT}(\delta_{\rm dmo})\times\sqrt{P_{yy,\rm hydro}/P_{\delta\delta,\rm hydro}})$. We find these alternatives result in larger covariances for the ratio $R$, and we therefore use the method described above since a tighter (or smaller) covariance estimate results in a more stringent test of our method.
@@ 4.1.1 Comparing Particle and Shell baryonification
For our tSZ profile painting method, we can interchangeably use the 3D and 2D approach as the difference is simply in what step the line-of-sight integration is done. [^16] This will no longer be the case if we used the approach of Aricò & Angulo (2024), who use the 3D locations of particles in the N-body simulation to generate a gas pressure field. In this case, there will be an associated error in using 2D projections of the 3D positions. In the first method, we paint profiles on a 3D grid (we do not use particle information when painting profiles) and then integrate along line-of-sight by collapsing the map along one direction. In the second, the integration is done within the definition of the projected, 2D profile, before any painting happens. Thus, the tSZ field does not suffer accuracy drops between using 3D fields and 2D fields.
@@
Once we have made the map for a given model, we take the fractional difference of the moments measured on the two maps from particle/shell baryonification, and then compute the average absolute ratio over all 32 models. Figure 3 shows this average difference for the moments of the density field. In practice, the offsets are within the 1-2% level, which is a negligible error. The error in the third moments increases at around the pixel-scale of $\approx 1\,{\rm Mpc}$, and the same for the fourth moments increases below $\lesssim 2\,{\rm Mpc}$. These differences are still within the cosmic variance-suppressed uncertainties of the measurement (see gray bands in Figure 8). We also show the $1\sigma$ spread of these absolute, fractional differences and find them to also be around the $1-2\%$ level for the scales mentioned above.
Figure 3: The absolute fractional difference in the baryonification predictions, averaged over 32 parameter choices, between the 2D shell method and the 3D particle method (see Section 2.2 for details on both methods). The dashed lines are the standard deviation of the absolute fractional differences. The 32 models are a Sobol sequence drawn from a prior defined in Table 1. The bias is within $2\%$ across all scales and increases for higher-order statistics for scales below $\lesssim 2\,{\rm Mpc}$. We deem this a negligible systematic; see text for more details.